OpenStudy (anonymous):

Help me check my work? You and your best friend want to backpack through Europe. You have done some research and discovered that it will cost $1600 for the plane ticket, rail pass, food, and lodging. You have already saved $900. If you invest this money in a savings account with a 2.3% interest rate compounded annually, how long will it take to earn enough money to go on the trip? Use the compound interest formula A= P (1 + i)n, where A is the accumulated amount, P is the principal, i is the interest rate per year, and n is the number of years. First you need to list the given variables

OpenStudy (anonymous):

like so: A=1600 P=$900 i=0.023 n=unknown Second you would set up your equation like so: 1600=900(1+0.023) After simplifying what's in the parenthesis the equation would look like this: 1600=(900+20.7)^n Next you would convert to a logarithm equation and solve like so: log (20.7) 1600=n log 1600/log 20.7=n n=2.4 years So it will take the two friends 2.4 years to save up the money for the trip if they invest it in a savings account with a 2.3% interest rate.

OpenStudy (welshfella):

1600 = 900*1.023^n 1.o23^n = 1600/900 = 1.7778 n log 1.023 = log 1.7778 n = log 1.7778 / log 1.023

OpenStudy (welshfella):

lines 7-9 in your solution are incorrect

OpenStudy (anonymous):

Alright thank you!

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

the 900 is not taken to the power n - Just the 1.023.